# Optimized Lattice-Structured Flexible EIT Sensor for Tactile Reconstruction and Classification

### v1

@@ IV-C HMI Application
V Tactile Classification
V-A Dataset and Network Training
To validate the superior spatio-temporal characteristics of our sensor, we conducted continuous motion detection experiments using various tactile interactions. Specifically, we selected tactile gestures with different movement trajectories as reference actions to assess the system’s performance. As shown in Fig. 10, the dataset includes 12 distinct interaction types such as no contact, finger press, four-finger scratch, fist press, finger double tap, palm pat, finger swipe up, finger swipe down, finger swipe left, finger swipe right, zoom-in and zoom-out. Each interaction type comprises 1000 samples, each containing 15 frames of voltage measurements. Then, we divided the 12,000 signals into a training set and a validation set in an 8:2 ratio, with 90% of the data used for model training and 10% for model validation and tuning. To fully evaluate the generalisation and robustness, 1,200 new signals were collected to be the test dataset, with each interaction type comprising 100 samples.
@@ V-A Dataset and Network Training
V-B Classification Results and Discussion

### v2

@@ V Human–Machine Interaction Demonstration
VI Tactile Classification: Temporal Trajectory Recognition
VI-A Dataset and Network Training
To validate the superior spatio-temporal characteristics of our sensor, we conducted continuous motion detection experiments using various tactile interactions. Specifically, we selected tactile gestures with different movement trajectories as reference actions to assess the system’s performance. As shown in Fig. 12, the dataset includes 12 distinct interaction types such as no contact, finger press, four-finger scratch, fist press, finger double tap, palm pat, finger swipe up, finger swipe down, finger swipe left, finger swipe right, zoom-in and zoom-out. Each interaction type comprises 1000 samples, each containing 15 frames of voltage measurements. Then, we divided the 12,000 signals into a training set and a validation set in an 8:2 ratio. To fully evaluate the generalisation and robustness, 1,200 new signals were collected to be the test dataset, with each interaction type comprising 100 samples.
@@ VI-A Dataset and Network Training
VI-B Classification Results and Discussion